OpenStudy (anonymous):

A number 15 is divided into 3 parts which are in Arithmetic Progression and the sum of their squares is 83 . What is the smallest part ?How to find it ? Plz help !

OpenStudy (anonymous):

let the first part be x.. thus second part= 15-x third part=15-x-x=15-2x x^2 + (15-x)^2 + (15-2x)^2= 83

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

but no real root for the equ if we do like this ?

OpenStudy (anonymous):

whaaaaaaattttt... wait i'll solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry the equation wont be formed like this wait ill tell u how to do it

OpenStudy (anonymous):

ohok

OpenStudy (anonymous):

15 is divided into three parts which are in arithmetic progression this means the numbers are consecutive let first part=x second part=x+1 third part=x+2 thus the equations are formed simultaneously 1st equation: x+(x+1)+(x+2)=15 2nd equation: \[x ^{2}+(x+1)^{2}+(x+2)^{2}=83\] solving 1st equation 3x+3=15 x=4 hopefully this is correct

OpenStudy (praxer):

is the sum of the numbers square is = 83.. Check it out once... @mohammad.arqum

OpenStudy (anonymous):

isn't it right @praxer ?

OpenStudy (praxer):

see, x was found to be 4 so the numbers will be 4,5 and 6 ..... right? but 4^2+5^2+6^2 = 77 not 83

OpenStudy (anonymous):

yes but answer is same i have done it in another way

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

same ques here too now !

OpenStudy (praxer):

...@mohammad.arqum it is 77 not 83... But according to the question it should be 83 ........right ?

OpenStudy (anonymous):

ok then the equation should become the second one \[x ^{2}+(x+1)^{2}+(x+2)^{2}=83\]

OpenStudy (anonymous):

but hw can we say that the common difference is 1 ? d can be any other number too right ?

OpenStudy (anonymous):

they are in arthemetic progressiom

OpenStudy (radar):

Arithmetic progression does not necessarily mean they have to be consecutive, while 4, 5, and 6 are a arithmetic progression, so is 2, 4, 6,

OpenStudy (anonymous):

mutherfckr

OpenStudy (praxer):

support @saravananr ... @mohammad.arqum follow @radar. A.P can have any common difference.

OpenStudy (anonymous):

@mohammad.arqum , why do u use foul language ?

OpenStudy (anonymous):

okee back to the question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so anyone know how to do em

OpenStudy (anonymous):

not me !

OpenStudy (anonymous):

wait

OpenStudy (radar):

I haven't resolved this problem, and would love to see it solved for my own education.

OpenStudy (anonymous):

yeah me too

OpenStudy (anonymous):

\[x+(x+n)+(x+2n)=15\] \[3x + 3n = 15\] \[n=5-x\] \[(x^2)+(x+n)^2+(x+2n)^2 = 83\] \[(x^2)+(x+5-x)^2+(x+10-2x)^2 = 83\] \[x^2+5^2+(10-x)^2=83\] that should work I think

OpenStudy (anonymous):

oh ! that's gr8 !

OpenStudy (anonymous):

\[x^2+25+100-20x+x^2-83=0\] \[2x^2-20x+42=0\] \[x^2-10x+21=0\]

OpenStudy (anonymous):

well i think it is best to solve the problem by the trail and error method The answer is 3,5, and 7 try it out yourselves

OpenStudy (anonymous):

Thanks a lot @AllTehMaffs

OpenStudy (radar):

Here are some arithmetic that sum up to 15 (other than 4, 5, and 6) 2, 5, and 8 1, 5, and 9 squares add up to 93

OpenStudy (radar):

How did I miss that one lol

OpenStudy (anonymous):

i think you r process iz better and logical than mine (y) @AllTehMaffs

OpenStudy (anonymous):

x=3 or 7 ! ?

OpenStudy (anonymous):

so whats the answer then

OpenStudy (anonymous):

It's 3 - even if the series is any fraction, 3*7>15

OpenStudy (anonymous):

~the common ratio

OpenStudy (anonymous):

the smallest part is three then what about remaining two parts

OpenStudy (radar):

Thanks @AllTehMaffs That seems to work.

OpenStudy (anonymous):

plug into n=5-x n=2 3,5,7

OpenStudy (anonymous):

yeah thats cool... :)

OpenStudy (radar):

3, 5, and 7 or 7, 5, 3

OpenStudy (anonymous):

thanks a lot @AllTehMaffs

OpenStudy (anonymous):

both are okeeee

OpenStudy (anonymous):

Oh! that's true! I forgot that n could be negative :P And welcome ^_^ Hooray maths!

OpenStudy (anonymous):

YAY!

OpenStudy (radar):

Glad this was worked, as I was beginning to get a headache.

OpenStudy (anonymous):

EVERYBODY TRY THIS ONE: http://openstudy.com/updates/52ac5794e4b0f72fdcf70305

OpenStudy (anonymous):

if n is -ve , it won't work here ! so ans is as already said , 3,5,7